# Structure-Aligned Protein Language Model

### v1

@@ 4.4 Mutation Effect Prediction
We evaluate our models on protein mutation effect prediction. Specifically, we consider three supervised tasks adopted in xTrimoPGLM (Chen et al., 2024): (1) Fluorescence: predicting the fluorescence intensity of green fluorescent protein mutants (Rao et al., 2019); (2) Fitness (GB1): predicting the binding fitness of GB1 following mutations at four specific positions; (3) Stability: predicting relative protease resistance as a proxy measurement for stability. For this task, evaluation is performed on one-mutation neighborhoods of the most promising proteins (Rao et al., 2019). We report performance using the Spearman correlation coefficient (denoted as Sp.). The setup is the same as in subsection 4.3, except that we fine-tune the backbone with a learning rate of $1\times 10^{-4}$.
@@
As shown in Table 2, SaESM2 demonstrates a clear advantage over other models in mutation effect prediction, particularly in Fitness (GB1) and Stability. We also observe that AMPLIFY models achieve performance comparable to ESM2 despite being almost half the size, highlighting the importance of preserving the natural distribution of proteins in the pre-training corpus.
4.5 Property Prediction
@@
We evaluate SaESM2 and SaAMPLIFY on a suite of downstream property prediction tasks (Xu et al., 2022; Dallago et al., 2021), which rely on structural information to some extent but are not direct structure prediction tasks. These include predictions of thermostability, metal ion binding, protein localization (DeepLoc), enzyme commission numbers (EC), gene ontology annotations (GO), and protein–protein interactions (HumanPPI). We follow the data splits and training protocols from SaProt (Su et al., 2024). While ESM2-based models are directly compatible with the SaProt codebase, AMPLIFY-based models are evaluated using a custom pipeline on the same splits, following the setup in subsection 4.4.
@@
SaESM2 outperforms ESM2 across $6$ out of $9$ downstream property prediction tasks, underscoring the advantages of integrating structural alignment into pLMs beyond structure-only downstream applications. Furthermore, SaESM2 achieves state-of-the-art performance among sequence-only pLMs, including structure-aligned baselines, on these $6$ tasks.
Table 3: Results on property prediction.
Method Thermostability HumanPPI Metal Bind EC GO (Fmax ($\uparrow$)) DeepLoc (Acc($\uparrow$))
Spearman ($\uparrow$) Acc($\uparrow$) Acc($\uparrow$) Fmax ($\uparrow$) MF BP CC Subcell. Binary
ESM2 $0.694$ $84.24$ $70.83$ 0.8640.864{0.864}0.864 $0.670$ $0.473$ $0.470$ $82.09$ $91.96$
ESM2-s $0.683$ $82.61$ $71.13$ $0.861$ 0.673 $0.479$ $0.458$ $82.85$ $92.28$
ISM $0.695$ $81.52$ $69.94$ 0.872 $0.666$ $0.471$ 0.4970.497{0.497}0.497 $82.63$ $92.28$
SaESM2 0.705 85.33 72.33 $0.861$ $0.669$ 0.488 $0.461$ 83.39 92.68
AMPLIFY $0.646$ $75.00$ $69.02$ $0.680$ $0.566$ $0.389$ $0.479$ $76.56$ $90.06$
SaAMPLIFY $0.675$ $71.11$ $70.83$ $0.773$ $0.586$ $0.412$ 0.512 $77.25$ $91.22$
4.6 Pseudo Perplexity
To measure the impact of the structure alignment on the pLMs sequence-level knowledge, we compute the pseudo perplexity distributions of ESM2, AMPLIFY, and their structure-aligned variants as defined in Section 1.2.2 of Lin et al. (2022) using the validation set from Fournier et al. (2024). This set includes proteins with experimental evidence from reference proteomes based on high-quality genomes across all three domains of life and is designed to better reflect the natural protein distribution.
Figure 2 reveals that our structure alignment does increases pseudo perplexity, indicating a trade-off in which structural integration slightly compromises sequence modeling. However, despite this, both SaESM2 650M and SaAMPLIFY 350M maintain competitive pseudo perplexity scores, suggesting that structure alignment largely preserves the original pLMs sequence-level knowledge.
Figure 2: Pseudo perplexity distributions on the validation set introduced by Fournier et al. (2024).
@@ 4.7 Ablation Studies
[…]
Table 4: Ablations on dual-task framework.
Contact Fluorescence Metal Bind
P@L/5 ($\uparrow$) Spearman ($\uparrow$) Acc($\uparrow$)
@@
w/o latent $53.7$ ($-12.0\%$) $0.689$ ($-0.9\%$) $69.5$ ($-3.8\%$)
w/o physical $59.1$ ($-3.1\%$) $0.691$ ($-0.6\%$) $71.0$ ($-1.8\%$)
w/o dual $51.4$ ($-15.7\%$) $0.686$ ($-1.3\%$) $67.1$ ($-7.2\%$)
ESM2 (baseline) $54.1$ ($-11.3\%$) $0.687$ ($-1.2\%$) $70.8$ ($-2.1\%$)
As shown in Table 4, removing any loss term leads to performance degradation across all three tasks, confirming the effectiveness of our dual-task framework. Notably, the w/o latent setting performs worse than w/o physical, suggesting that the latent-level task contributes more significantly to the considered downstream tasks than the physical-level task. This supports our motivation that the physical-level task acts primarily as a structural constraint rather than a dominant learning signal.
@@
Contact Fluorescence Metal Bind
P@L/5 ($\uparrow$) Spearman ($\uparrow$) Acc($\uparrow$)
@@
loss-large $60.6$ (--0.7%) $0.693$ (--0.3%) $71.3$ (--1.4%)
loss-small $59.4$ (--2.6%) $0.691$ (--0.6%) $71.0$ (--1.8%)
full $60.3$ (--1.1%) $0.690$ (--0.7%) $71.1$ (--1.7%)
ESM2 (baseline) $54.1$ (--11.3%) $0.687$ (--1.2%) $70.8$ (--2.1%)
As shown in Table 5, alternative selection strategies lead to decreased performance across all tasks, demonstrating the effectiveness of our residue loss selection module. While beneficial, its impact is less significant than that of the dual-task framework, likely due to the already high quality of the protein structures used and the extensive pre-training of base pLMs. We further visualize the validation loss curves for different loss selection strategies in §D, which further supports the superior effectiveness of our strategy.
@@
Contact Fluorescence Metal Bind
P@L/5 ($\uparrow$) Spearman ($\uparrow$) Acc($\uparrow$)
@@
AF2 $48.4$ (--20.7%) 0.695 (--0.0%) $69.0$ (--4.6%)
ESM2 (baseline) $54.1$ (--11.3%) $0.687$ (--1.2%) $70.8$ (--2.1%)
As shown in Table 6, aligning to GearNet embeddings outperforms AF2 alignment on both Contact Prediction and Metal Bind tasks. Notably, AF2 alignment even underperforms the baseline ESM2 model without structural alignment. This observation is consistent with the findings of Hu et al. (2022), which suggest that embeddings from the AF2 are not well-suited for such downstream tasks.
@@
Contact Fluorescence Metal Bind
P@L/5 ($\uparrow$) Spearman ($\uparrow$) Acc($\uparrow$)
SaESM2 (foldseek) 61.061.0{61.0}61.0 0.695 72.3
protoken 60.860.8{60.8}60.8 (--0.3%) 0.695 (+++0.0%) $71.9$ (--0.6%)
aido 61.9 ($+$1.5%) 0.695 ($+$0.0%) $70.5$ (--2.5%)
ESM2 $54.1$ (--11.3%) $0.687$ (--1.2%) $70.8$ (--2.1%)
@@ 5 Conclusion
In this work, we propose to enrich sequence-only pLMs with structural knowledge. We incorporate structural insights from pre-trained pGNNs into pLMs via a latent-level task, aligning residue representations across models. To infuse intra-protein structural knowledge, we introduce a physical-level task that trains pLMs to predict structural tokens. Additionally, we propose a residue loss selection module that identifies and emphasizes challenging yet reliable residue losses to guide learning. We validate our structure alignment approach on two pLMs, ESM2 and AMPLIFY, demonstrating improved performance across diverse downstream tasks. These results suggest that structure alignment could become an indispensable component for future pLMs.
Appendix A Terminology
Appendix B Related Work
Structure Language Models
There are two main types of structure language models. The first requires explicit structural input, such as structure tokens (Su et al., 2024; Heinzinger et al., 2024; Li et al., 2024) or torsion angles (Frolova et al., 2024). However, these models depend on potentially unreliable or inaccurate structural data, and protein structure databases like the PDB are much smaller than sequence-only databases. Additionally, many proteins lack a well-defined, rigid structure. The second type only requires protein sequences as input and integrates structural insights during pre-training. For example, Zhang et al. (2024b) introduces a physical-level task for fold prediction, though it is somewhat coarse. Sun and Shen (2024) proposes several physical-level tasks, including secondary structure and distance map predictions, to incorporate structural knowledge into the pLM, while Ouyang-Zhang et al. (2024) focuses on structure token prediction. Penaherrera and Koes (2024) uses a similar contrastive learning loss, but limits its focus to masked residues and does not utilize advanced pre-trained GNN models. Finally, AlphaFold2 (Jumper et al., 2021) and ESMFold (Lin et al., 2023b) use sequence encoders, namely Evoformer and ESM2, followed by structure prediction modules. However, their focus is on structure prediction, and AlphaFold2 embeddings have been shown to be less effective than ESM2 embeddings for downstream tasks (Hu et al., 2022).
While recent studies have explored how to incorporate knowledge from pre-trained pLMs into pGNNs (Zheng and Li, 2024; Chen et al., 2023; Robinson et al., 2023), their focus is on improving pGNNs rather than pLMs, and no prior work has explored integrating structural insights from pre-trained pGNNs into pLMs. Our work bridges this gap by introducing the latent-level task, thereby enriching the pLMs with comprehensive structural insights from pre-trained pGNNs.
Data Selection
@@ Appendix B Related Work
Appendix C Residue Embedding Visualization

### v2

@@ 4.3.3 Supervised Mutation Effect Prediction
We evaluate our models on protein mutation effect prediction. Specifically, we consider two supervised tasks adopted in xTrimoPGLM: (1) Fitness (GB1): predicting the binding fitness of GB1 following mutations at four specific positions; (2) Stability: predicting relative protease resistance as a proxy measurement for stability. For this task, evaluation is performed on one-mutation neighborhoods of the most promising proteins (Rao et al., 2019). We report performance using the Spearman correlation coefficient, while the setup is the same as in §4.3.1, except that we also fine-tune the backbone with a learning rate of $1\times 10^{-4}$.
@@
As shown in Table 1, SaESM2 demonstrates a clear advantage over other models in supervised mutation effect prediction in Stability prediction  [^3] Note that standard deviation of the downstream evaluation pipeline for Stability prediction is very high (see Table 9 in Appendix) and a statistically competitive performance for Fitness prediction (GB1).
Additional results, across model sizes and downstream tasks can be found in Appendix C.1, providing additional insights into how the method scales. In the same section, we also report our statistically inconclusive results on Fluorescence prediction under protein mutations.
4.4 Zero-Shot Deep Mutational Scanning
@@
We evaluate all our models on zero-shot deep mutational scanning (DMS), which is related to the supervised mutation effect prediction tasks in §4.3.3 but evaluated in a zero-shot setting. Here, we use the model’s masked language modeling head to score mutations, comparing the log-likelihood of the mutated amino acid to the wild-type. We use the ProteinGym (Notin et al., 2023) DMS substitution benchmark, which compares predicted scores to experimental fitness scores for 217 assays. We used public figures for the ESM2, ESM-C and ESM3 families of models.
@@
As shown in Figure 3, the structure-aligned models significantly outperform their unaligned counterparts across model families and sizes.
This consistent improvement in zero-shot fitness prediction is a strong indicator of an increase in biophysical understanding.
Protein language models rely on sequence co-evolution to predict fitness, that is, essentially predicting fitness based on "what amino acid is common at this position?". In contrast, DMS assays measure experimental fitness, which is often dominated by protein stability. Our structure-aligned models overcome this limitation. By infusing inter-protein and intra-protein structural knowledge, they constrain the MLM head to favor mutations that are not just sequentially plausible but also structurally sound. A mutation that would destabilize the protein’s fold is now correctly assigned a lower probability, leading to a stronger correlation with the experimental fitness data. Additional details about ProteinGym, including violin plots for all model and a head-to-head comparison are provided in Appendix C.3.
Figure 3: DMS average Spearman correlation scaling against model size for different families of models. Structure-aligned models, identified by the prefix Sa-*, consistently outperform their baseline models. SaESM2 650M and 150M are now on the Pareto front of performance against model size defined by ESM-C and ESM3.
4.5 Pseudo-perplexity
To measure the impact of the structure alignment on the pLMs’ sequence-level knowledge, we compute the pseudo-perplexity distributions of ESM2, AMPLIFY, and their structure-aligned variants as defined in Section 1.2.2 of Lin et al. (2022) using the validation set from Fournier et al. (2024). This set includes proteins with experimental evidence from reference proteomes based on high-quality genomes across all three domains of life and is designed to reflect accurately the natural protein distribution.
Figure 4 reveals that our structure alignment does increase pseudo-perplexity, indicating a trade-off in which structural integration slightly compromises sequence modeling. However, despite this, both SaESM2 650M and SaAMPLIFY 350M maintain competitive pseudo-perplexity scores, suggesting that structure alignment largely preserves the original pLMs sequence-level knowledge.
Table 2: Mean Perplexity (PPL) on the test set with 95% Confidence Intervals. T-tests compare the base model (ESM2, AMPLIFY) against its corresponding Sa-variant (SaESM2, SaAMPLIFY).
Family Model Mean PPL [95% CI] t-statistic p-value
ESM2 ESM2 (650M) 5.89 [5.81, 5.96] -8.65 $5.50\times 10^{-18}$
SaESM2 (650M) 6.38 [6.29, 6.46]
AMPLIFY AMPLIFY (350M) 4.58 [4.50, 4.65] -9.30 $1.55\times 10^{-20}$
SaAMPLIFY (350M) 5.10 [5.02, 5.18]
Figure 4: Pseudo-perplexity distributions on the validation set introduced by Fournier et al. (2024).
@@ 4.6 Ablation Studies
Our default setup employs a weighted combination of three losses: masked language modeling, latent-level, and physical-level, with weights $\left(1,\,0.5,\,0.5\right)$, respectively. To assess the impact of each component, we experiment with the following configurations:
  • w/o latent: Remove the latent-level loss, using weights $\left(1,\,0,\,0.5\right)$.
  • w/o physical: Remove the physical-level loss, using weights $\left(1,\,0.5,\,0\right)$.
  • w/o dual: Exclude both auxiliary losses, i.e. MLM fine-tuning on PDB, using weights $\left(1,\,0,\,0\right)$.
Table 3: Ablations on dual‑task framework.
Contact on the trRosetta split Fluorescence Metal Bind
P@L/5 ($\uparrow$) Spearman ($\uparrow$) Acc% ($\uparrow$)
@@
w/o latent $53.7$ ($12.0\%$) $0.689$ ($0.9\%$) $69.5$ ($3.8\%$)
w/o physical $59.1$ ($3.1\%$) $0.691$ ($0.6\%$) $71.0$ ($1.8\%$)
w/o dual $51.4$ ($15.7\%$) $0.686$ ($1.3\%$) $67.1$ ($7.2\%$)
ESM2 (baseline) $54.1$ ($11.3\%$) $0.687$ ($1.2\%$) $70.8$ ($2.1\%$)
As shown in Table 3, removing any loss term leads to performance degradation across all three tasks, confirming the effectiveness of our dual-task framework. Notably, the w/o latent setting performs worse than w/o physical, suggesting that the latent-level task contributes more significantly to the considered downstream tasks than the physical-level task. This supports our motivation that the physical-level task acts primarily as a structural constraint rather than as a dominant learning signal.
@@
Contact on the trRosetta split Fluorescence Metal Bind
P@L/5 ($\uparrow$) Spearman ($\uparrow$) Acc% ($\uparrow$)
@@
loss-large $60.6$ (-0.7%) $0.693$ (-0.3%) $71.3$ (-1.4%)
loss-small $59.4$ (-2.6%) $0.691$ (-0.6%) $71.0$ (-1.8%)
full $60.3$ (-1.1%) $0.690$ (-0.7%) $71.1$ (-1.7%)
ESM2 (baseline) $54.1$ (-11.3%) $0.687$ (-1.2%) $70.8$ (-2.1%)
As shown in Table 4, alternative selection strategies led to decreased performance across all tasks, demonstrating the effectiveness of our residue loss selection module. While beneficial, its impact is less significant than that of the dual-task framework, likely due to the already high quality of the protein structures used and the extensive pre-training of base pLMs. We further visualize the validation loss curves for different loss selection strategies in §D, which further supports the superior effectiveness of our strategy.
@@
Contact on the trRosetta split Fluorescence Metal Bind
P@L/5 ($\uparrow$) Spearman ($\uparrow$) Acc% ($\uparrow$)
@@
AF2 $48.4$ (-20.7%) 0.695 (-0.0%) $69.0$ (-4.6%)
ESM2 (baseline) $54.1$ (-11.3%) $0.687$ (-1.2%) $70.8$ (-2.1%)
As shown in Table 5, aligning with GearNet embeddings outperforms aligning with AlphaFold2 embeddings on both Contact Prediction (trRosetta split) and Metal Bind tasks. We also observed a degradation of our method when aligning to AF2 embeddings compared to the baseline ESM2 model without structural alignment. This observation is consistent with the findings of Hu et al. (2022), which suggest that embeddings from the AF2 may not be well-suited for some downstream tasks.
@@
Contact on the trRosetta split Fluorescence Metal Bind
P@L/5 ($\uparrow$) Spearman ($\uparrow$) Acc% ($\uparrow$)
SaESM2 (foldseek) 61.0{61.0} 0.695 72.3
protoken 60.8{60.8} (-0.3%) 0.695 (++0.0%) $71.9$ (-0.6%)
aido 61.9 (++1.5%) 0.695 (++0.0%) $70.5$ (-2.5%)
ESM2 $54.1$ (-11.3%) $0.687$ (-1.2%) $70.8$ (-2.1%)
@@ 5 Conclusion
In this work, we propose to enrich sequence-only pLMs with structural knowledge. We incorporate structural insights from pre-trained pGNNs into pLMs via a latent-level task, aligning residue representations across models. To infuse intra-protein structural knowledge, we introduce a physical-level task that trains pLMs to predict structural tokens. Additionally, we propose a residue loss selection module that identifies and emphasizes challenging yet reliable residue losses to guide learning. We validate our structure alignment approach on two pLMs, ESM2 and AMPLIFY, demonstrating improved performance across diverse downstream tasks, from structure prediction to supervised and unsupervised mutation effect prediction tasks. These results suggest that structure alignment could become an indispensable component for future pLMs.
Appendix A Related Work
A.1 Structure Language Models
There are two main types of structure language models. The first requires explicit structural input, such as structure tokens (Su et al., 2024; Heinzinger et al., 2024; Li et al., 2024) or torsion angles (Frolova et al., 2024) or geometric graphs (Hartout et al., 2025). However, these models depend on potentially unreliable or inaccurate structural data (protein structures are generally modeled at cryogenic temperatures and fail to take into account the full conformational landscape) and protein structure databases like the PDB are much smaller than sequence-only databases. Additionally, many proteins lack a well-defined, rigid structure, having disordered domains. All these approaches are not directly comparable to ours as their models requires explicit structural inputs, whereas SaESM operates purely on sequences, a property we believe is important.
[…]
While recent studies have explored how to incorporate knowledge from pre-trained pLMs into pGNNs (Zheng and Li, 2024; Chen et al., 2023; Robinson et al., 2023), their focus is on improving pGNNs rather than pLMs, and no prior work has explored integrating structural insights from pre-trained pGNNs into pLMs. Our work bridges this gap by introducing the latent-level task, thereby enriching the pLMs with comprehensive structural insights from pre-trained pGNNs. Finally, the idea of distilling structural information via some form of contrastive learning between sequences is not new, with (Bepler and Berger, 2021) directly predicting contact inside a protein while simultaneously contrasting SCOP (Chandonia et al., 2017) information between protein pairs, with a language modeling trunk.
A.2 Data Selection
@@
Appendix B Residue Embedding Visualization
@@
Table 7: Quantitative evaluation of embedding separability. We report the Silhouette Score (measure of cluster cohesion/separation between -1 and 1, higher is better) and k-NN Classification Accuracy ($k=20$) for residue type, grouped residue properties as in Figure 6, and Secondary Structure (Q3) labels. Bold values indicate the best performance within each model family.
Silhouette Score k-NN Accuracy ($k=20$)
Model Amino Acid Grouped AA Sec. Str. (Q3) Amino Acid Grouped AA Sec. Str. (Q3)
ESM2 0.023 0.005 -0.001 0.981 0.983 0.772
SaESM2 0.030 0.010 0.012 0.964 0.967 0.861
AMPLIFY 0.053 0.029 -0.007 0.929 0.946 0.644
SaAMPLIFY 0.058 0.027 0.009 0.956 0.964 0.798
@@
Appendix C Additional Results and Scaling Analysis
We provide in this section additional results and visualizations on downstream property prediction benchmarks. As shown in the figures, our method scales favorably on tasks where scaling model size helps. On all tasks, considering confidence intervals, our method’s performance is either comparable or higher than the base model performance. In Figure 11, we provide full violin plots for all our ProteinGym evaluations.
C.1 Full Results on downstream property prediction: xTrimoPGLm
In this section, we report the full results, across model sizes on every evaluation from xTrimoPGLM we conducted. In Figure 7, we show that, after averaging scores on xTrimoPGLM evaluations, struture-aligned models outperform their baseline, with the largest gaps observed starting in the 100M parameters range. In Figure 8, we plot for every task all of our results.
Considering structure based tasks that are not contact map (already discussed in Figure 2), we find consistent improvements in secondary structure prediction across all model sizes. For fold prediction for all SaESM2 models up to the 150M parameters size have significantly higher accuracy. For the 650M model, confidence intervals overlap while ESM-S, for which the full backbone was trained in remote homology detection, remains best. SaAMPLIFY models always outperform their baseline.
In term of supervised mutation effect prediction, we are only able to report results for all model size for stability prediction, where SaESM2 has higher accuracy at both the 150M and 650M model sizes.
Finally, for the three last tasks: peptide-HLA MHC affinity, TCR-pMHC affinity and enzyme catalytic efficiency, we cannot report significant results.
Figure 7: Average model performance compared with model size for different families of models over xTrimoPGLM. Gaps between models would widen if downstream evaluations with fully overlapping confidence intervals between all methods and baselines were removed.
Figure 8: Model performance compared with model size for different families of models over every xTrimoPGLM task.
C.2 Full Results on downstream property prediction: SaProt
In this section, we report the full results, across model sizes on every evaluation from SaProt we conducted. As shown in Figure 9, no improvement can be observed with structure alignment. Only a general scaling of performance with model size is visible.
The full results are displayed in Figure 10, with figures at the largest model scales in Table 8. No model outperforms significantly any of the others. For most of the downstream evaluations, confidence intervals bootstrapped from the test set are completely overlapped. We offer three possible interpretations, both for SaProt and for inconclusive xTrimoPGLM tasks: (i) these downstream tasks either do not exhibit or do not benefit from transfer learning of a foundation model (ii) the training data for these benchmarks tasks is fairly noisy (iii) the test split size for these tasks is too small compared to their noise levels, yielding large confidence intervals.
Table 8: Results on functional property prediction. Values are Metric [95% Confidence Interval]. Within each model family (ESM2-based and AMPLIFY-based), the best-performing model is bolded. Models in italics have a mean score that falls within the 95% CI of the best model in their family. SaESM2 is within confidence intervals for all tasks where it’s not best, apart from GO Cellular Component and Human PPI. Confidence intervals for models (†) and tasks (†) are statistical upper bounds.
EC GO (BP) GO (CC) GO (MF) Thermostability
Model Fmax ($\uparrow$) Fmax ($\uparrow$) Fmax ($\uparrow$) Fmax ($\uparrow$) Sp. ($\uparrow$)
ESM2 0.855 [0.841 2900.841\,290, 0.869 4380.869\,438] 0.477 [0.467 0300.467\,030, 0.486 6570.486\,657] 0.484 [0.473 5350.473\,535, 0.493 2220.493\,222] 0.672 [0.660 7250.660\,725, 0.681 2470.681\,247] 0.712 [0.572 7500.572\,750, 0.764 5350.764\,535]
ESM-S† 0.861 [0.844 4940.844\,494, 0.877 5430.877\,543] 0.479 [0.462 2940.462\,294, 0.496 1280.496\,128] 0.458 [0.440 8100.440\,810, 0.474 5520.474\,552] 0.673 [0.657 1600.657\,160, 0.688 7160.688\,716] 0.683 [0.658 1530.658\,153, 0.707 5430.707\,543]
ISM† 0.872 [0.855 6680.855\,668, 0.887 5640.887\,564] 0.471 [0.454 3740.454\,374, 0.488 1160.488\,116] 0.497 [0.479 7720.479\,772, 0.513 4950.513\,495] 0.666 [0.649 7600.649\,760, 0.681 6460.681\,646] 0.695 [0.670 6080.670\,608, 0.719 2310.719\,231]
S-PLM 0.878 [0.866 3440.866\,344, 0.891 7540.891\,754] 0.480 [0.471 7210.471\,721, 0.490 5110.490\,511] 0.445 [0.434 7680.434\,768, 0.454 8460.454\,846] 0.671 [0.659 8880.659\,888, 0.681 7760.681\,776] 0.704 [0.590 1800.590\,180, 0.766 4690.766\,469]
SaESM2 (ours) 0.868 [0.855 1560.855\,156, 0.881 7170.881\,717] 0.479 [0.469 8080.469\,808, 0.488 5830.488\,583] 0.462 [0.451 6270.451\,627, 0.472 6590.472\,659] 0.663 [0.652 5580.652\,558, 0.673 7940.673\,794] 0.693 [0.570 1720.570\,172, 0.755 9000.755\,900]
AMPLIFY 0.501 [0.480 0330.480\,033, 0.524 6310.524\,631] 0.271 [0.263 2640.263\,264, 0.279 0760.279\,076] 0.322 [0.311 0620.311\,062, 0.332 2240.332\,224] 0.378 [0.366 2380.366\,238, 0.392 5380.392\,538] 0.614 [0.430 4130.430\,413, 0.640 2140.640\,214]
SaAMPLIFY (ours) 0.486 [0.463 9460.463\,946, 0.508 3290.508\,329] 0.257 [0.249 5650.249\,565, 0.265 9230.265\,923] 0.348 [0.341 7650.341\,765, 0.358 1000.358\,100] 0.389 [0.376 3210.376\,321, 0.400 9100.400\,910] 0.596 [0.420 0730.420\,073, 0.641 0920.641\,092]
DeepLoc (Subcell.) DeepLoc (Binary) HumanPPI Metal Bind
Model Acc ($\uparrow$) Acc ($\uparrow$) Acc ($\uparrow$) Acc ($\uparrow$)
ESM2 0.839 [0.824 9070.824\,907, 0.851 9510.851\,951] 0.931 [0.918 9260.918\,926, 0.942 1920.942\,192] 0.783 [0.722 3510.722\,351, 0.836 2440.836\,244] 0.705 [0.670 7080.670\,708, 0.739 6910.739\,691]
ESM-S† 0.828 [0.814 3920.814\,392, 0.842 1090.842\,109] 0.934 [0.922 6630.922\,663, 0.945 3710.945\,371] 0.826 [0.770 7140.770\,714, 0.875 1080.875\,108] 0.711 [0.676 8670.676\,867, 0.745 4700.745\,470]
ISM† 0.826 [0.812 1410.812\,141, 0.839 9980.839\,998] 0.923 [0.910 2380.910\,238, 0.934 7420.934\,742] 0.815 [0.758 5170.758\,517, 0.865 5000.865\,500] 0.699 [0.664 5550.664\,555, 0.733 9040.733\,904]
S-PLM 0.847 [0.833 9390.833\,939, 0.860 1020.860\,102] 0.930 [0.917 0510.917\,051, 0.941 8200.941\,820] 0.853 [0.798 9130.798\,913, 0.902 1740.902\,174] 0.696 [0.660 7140.660\,714, 0.730 6550.730\,655]
SaESM2 (ours) 0.840 [0.826 4110.826\,411, 0.853 3550.853\,355] 0.933 [0.920 7930.920\,793, 0.943 7830.943\,783] 0.777 [0.716 3800.716\,380, 0.831 3100.831\,310] 0.759 [0.726 4180.726\,418, 0.789 9990.789\,999]
AMPLIFY 0.689 [0.672 0110.672\,011, 0.706 3970.706\,397] 0.861 [0.844 8860.844\,886, 0.876 7490.876\,749] 0.690 [0.622 6650.622\,665, 0.755 6680.755\,668] 0.621 [0.583 6540.583\,654, 0.656 5990.656\,599]
SaAMPLIFY (ours) 0.674 [0.656 1580.656\,158, 0.690 9880.690\,988] 0.881 [0.865 4940.865\,494, 0.895 3360.895\,336] 0.734 [0.669 1200.669\,120, 0.796 3720.796\,372] 0.601 [0.563 9740.563\,974, 0.637 4710.637\,471]
Figure 9: Average model performance compared with model size for different families of models over SaProt. The comparatively larger gap between AMPLIFY and ESM based models between the SaProt and xTrimoPGLM evaluation is probably the consequence of observed hyper-parameter sensitivity in the former evaluation pipeline.
Figure 10: Model performance compared with model size for different families of models over every xTrimoPGLM task. Confidence intervals for models or tasks with $\dagger$} are statistical upper-bounds on confidence intervals. For other tasks and models, confidence intervals are computed via bootstrapping on the test set.
C.3 Additional Results on ProteinGym
In Figure 11, we present full violin plots for all models involved in our comparison, while Figure 12 compares the performance of our models on every assay for the original model and the structure aligned one. For smaller size models, our method seems to reduce the number of assays where our fitness predictions are anti-correlated with the ground truth. For every model family and size, the improvements seem to be due in equal part to: (i) a slight improvement on average on most of the assays (ii) some assays where Spearman correlations are substantially improved. A closer inspection of these assays over the three standard ProteinGym assay metadata information (Taxon, MSA depth and Selection type) reveals no correlation between the assay type and the improvement from our method.
Overall, our aligned SaESM2 outperforms ESM2 3B and 15B (older and bigger models) and both ESM-C 300M and 600M (models of the new generation). ESM3 still remains better on ProteinGym, although he is significantly larger and fully multimodal.
Figure 11: Violin plots of the distribution of assay spearman correlations for all models evaluated on ProteinGym. The solid black line at 0 represents the expected correlation of a random model.
Figure 12: Head-to-head comparison of model performance on ProteinGym. Each point represents a single deep mutational scanning (DMS) assay. (Top) SaESM2 (y-axis) vs. ESM2 (x-axis). (Bottom) SaAMPLIFY (y-axis) vs. AMPLIFY (x-axis). Points above the $y=x$ diagonal (dashed line) indicate an improvement over the assay from our structure alignment.
C.4 Stability of our post-training method: analysis of the downstream performance over 3 seeds.
Table 9: Robustness of the post-training procedure. We compare the standard deviation of the final test metric across 3 independent post-training seeds (Std. Dev. (3 Seeds)) against the 95% confidence interval (CI) size computed via bootstrapping on the test set (Test Set CI Size). The low standard deviation across seeds demonstrates the high stability of our alignment method.
Task Std. Dev. (3 Seeds) Test Set CI Size
Contact Prediction CASP16 (Long Range P@L) 0.0031 0.0204
Enzyme Catalytic Efficiency (Pearson Correlation) 0.0022 0.0705
Fitness Prediction (Spearman Correlation) 0.0027 0.0108
Fluorescence Prediction (Spearman Correlation) 0.0084 0.0160
Fold Prediction (Accuracy) 0.0026 0.0309
Peptide-HLA MHC Affinity (AUC) 0.0027 0.0066
Secondary Structure Prediction (Accuracy) 0.0013 0.0032
Stability Prediction (Accuracy) 0.0766 0.0185
TCR-pMHC Affinity (AUC) 0.0022 0.0154
To validate the stability of our structure-alignment procedure, we performed post-training using 3 different random seeds and measured the standard deviation of the final test metric. In Table 9, we compare this post-training variance against the inherent uncertainty of the benchmarks themselves, represented by the 95% confidence interval (CI) size derived from bootstrapping the test set.For 8 out of the 9 tasks, the standard deviation from our post-training seeds is substantially smaller than the test set CI, often by an order of magnitude (e.g., $0.0027$ vs. $0.0108$ for Fitness Prediction). This result is crucial, as it indicates that our method is highly robust and that the observed variance in benchmark scores is dominated by the test set’s composition, not by the stochasticity of our alignment process. We note that Stability Prediction is an exception, showing higher seed variance than test set uncertainty. After further investigation, this sensitivity is mainly due to the fine-tuning pipeline, as we found our results on Stability Prediction to be particularly sensitive to hyper-parameter choices.